OpenStudy (anonymous):

i really need help on this: You buy a game for $40, and the sales tax is $2. What percent of the game price are you paying in sales tax?

OpenStudy (rock_mit182):

x =percent of the game price you're paying in sales tax

OpenStudy (rock_mit182):

40/2=100/x -->\[x =2 *100/40\]